### v1

# Kinematify: Open-Vocabulary Synthesis of High-DoF Articulated Objects

Abstract
A deep understanding of kinematic structures and movable components is essential for enabling robots to manipulate objects and model their own articulated forms. Such understanding is captured through articulated objects, which are essential for tasks such as physical simulation, motion planning, and policy learning. However, creating these models, particularly for complex systems like robots or objects with high degrees of freedom (DoF), remains a significant challenge. Existing methods typically rely on motion sequences or strong assumptions from hand-curated datasets, which hinders scalability. In this paper, we introduce Kinematify, an automated framework that synthesizes articulated objects directly from arbitrary RGB images or text prompts. Our method addresses two core challenges: (i) inferring kinematic topologies for high-DoF objects and (ii) estimating joint parameters from static geometry. To achieve this, we combine MCTS search for structural inference with geometry-driven optimization for joint reasoning, producing physically consistent and functionally valid descriptions. We evaluate Kinematify on diverse inputs from both synthetic and real-world environments, demonstrating improvements in registration and kinematic topology accuracy over prior work.
22footnotetext: Project lead: Qixuan Zhang (zhangqx@deemos.com).11footnotetext: Corresponding authors: Jingyi Yu (yujingyi@shanghaitech.edu.cn), Lan Xu (xulan1@shanghaitech.edu.cn).
I Introduction
Enabling robots to effectively interact with and manipulate objects, as well as to model their own articulated structures for self-perception and adaptation, requires a deep understanding of kinematic topologies and joint parameters. Articulated robot descriptions capture this understanding by encoding geometry, kinematic dependencies, and dynamic constraints in standard formats like the Unified Robot Description Format (URDF) [1]. These descriptions are essential for core robotic tasks such as manipulation, locomotion, and policy learning. However, customizing such descriptions for articulated objects remains a significant challenge, demanding substantial manual effort, especially for high degrees of freedom (DoF) systems like humanoids, quadrupeds, and arms. This difficulty arises from the labor intensive processes of part-aware 3D modeling, resolving intricate kinematic dependencies, and inferring precise joint parameters [2].
While recent advances in part-aware 3D generation [3, 4, 5, 6, 7] now enable the on-demand creation of high-quality segmented meshes from text or image prompts, the bottleneck of kinematics inference remains. This challenge has driven robotics researchers to explore high-DoF articulated objects generation approaches.
Prior work has followed two main directions. Geometry-first approaches infer parts and joints from dense 4D sequences or multi-scan data, which achieve high fidelity but rely on controlled capture settings [8, 9]. Program-synthesis pipelines, in contrast, predict executable descriptions directly from visual inputs [10, 11, 12]. While effective, these systems mainly target everyday objects such as laptops, bottles, and drawers, which typically contain only a few moving parts and relatively simple kinematic dependencies. In the context of robot self-modelling, related work such as AutoURDF [13] extends this idea to robots, recovering topology and joint types from point-cloud sequences. However, it presumes motion data and is largely limited to serial-chain structures, whereas high-DoF objects often exhibit multi-branched linkages.
We introduce Kinematify, a pipeline that synthesizes articulated robot descriptions from RGB images or texts. Our method first generates the segmented mesh with a 3D foundation model, then constructs a contact graph and infers a kinematic tree using an MCTS [14, 15, 16] objective that balances hierarchy and structural regularity. Subsequently, it estimates joint parameters via DW-CAVL, a novel Distance-Weighted Contact-Aware Virtual Linkage optimization approach. This approach preserves near-contact regions while penalizing collisions under virtual motion. The resulting description is exported to URDF and is readily convertible to formats like MJCF [17] or USD [18]. Our contributions are:
  • An automated pipeline for articulated object synthesis. Kinematify generates physics-aware, high-DoF articulated objects directly from arbitrary RGB images or text prompts, without requiring motion data or pre-defined articulation priors.
  • MCTS-based kinematic tree inference. We propose a search objective that encodes structural priors like hierarchy and regularity to resolve ambiguous attachments for complex, high-DoF articulated objects.
  • A geometry-driven joint parameter estimation method. Our DW-CAVL algorithm accurately infers revolute and prismatic joint parameters from static geometry alone by optimizing an SDF-based, contact-aware objective under virtual motions.
II Related Work
II-A 3D Articulated Object Modelling
A significant body of work focuses on reconstructing the kinematic structure of everyday objects from visual data. The most common paradigm leverages motion to reveal articulation [9, 8, 19, 20, 13]. By observing an object over time, these methods can directly infer which parts move together and identify the axes of rotation or translation. For example, MultiBodySync [9] registers multiple 3D scans of an object in different states, using spectral synchronization to jointly solve for part segmentation and motion. Similarly, ReArt [8] fits a rearticulable model to a 4D point cloud sequence by jointly optimizing segmentation, topology, and joint parameters. These methods achieve high fidelity but depend critically on the availability of multi-view or temporal data, which requires a controlled capture setup. Another trend frames articulation modelling as a program synthesis problem [12, 10, 11, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. URDFormer [12] trains a transformer to predict a URDF from a single image, relying on a large-scale synthetic dataset of image-URDF pairs. Real2Code [10] and Articulate-Anything [11] leverage large language models (LLMs) to generate code-based representations of articulated objects, with the latter using a reinforcement learning loop to refine the model through simulation feedback. While powerful in their open-vocabulary capabilities, these methods often produce functionally plausible rather than geometrically precise models and struggle with the complex topologies common in high-DoF objects.
II-B Robot Self-Modeling
Distinct from general everyday object modeling, robot self-modeling is the online process by which a robot autonomously discovers its own body plan [34, 35, 36, 13, 37, 38]. This is typically achieved by correlating motor actions with sensory feedback. The foundational concept of task-agnostic self-modeling [34] involves a robot performing random motions and building a self-representation from the resulting data. This has been realized in various ways. Ledezma et al. [36] used IMU sensors on each link, applying machine learning to the sensor data to explicitly solve for the robot’s topology and kinematic parameters. These methods are powerful but require an embodied agent with access to its own motor and sensory signals. AutoURDF [13] represents a purely visual approach to this problem. It operates on time-series point cloud frames of a robot in motion, but without access to the underlying motor commands. By tracking the 6-DoF transformations of point clusters, it segments moving parts, infers the kinematic tree using a minimum spanning tree, and estimates joint parameters. It demonstrated superior performance in topology inference for serial-chain objects.
In contrast to robot self-modelling methods that require an observation of the robot, Kinematify operates in a zero-shot physical context, inferring the kinematics topology from the image or text alone. These open vocabulary settings make our approach applicable much earlier in the design pipeline, before a physical prototype is built or motion data is available. To the best of our knowledge, Kinematify is the first method to construct high-DoF, kinematics-aware articulated objects directly from unlabelled RGB images or texts.
III Kinematify
To overcome the limitations of prior self-modelling methods, we introduce Kinematify, a zero-shot framework that constructs high-DoF, kinematics-aware articulated objects directly from unlabelled RGB images or textual descriptions.
III-A Preliminaries
Assemblies and parts. An assembly $\mathcal{A}$ is a finite set of parts $P=\{P_{i}\}_{i=1}^{N}$. Part $P_{i}$ has a triangulated surface mesh $M_{i}=(V_{i},F_{i})$ with vertices $V_{i}\subset\mathbb{R}^{3}$ and triangular faces $F_{i}\subset\{1,\dots,|V_{i}|\}^{3}$. Each part stores a world transform $T_{i}\in\mathrm{SE}(3)$, an intrinsic rotation $R_{i}\in\mathrm{SO}(3)$ used for alignment, a centroid $c_{i}\in\mathbb{R}^{3}$, a robust volume $v_{i}>0$, and axis-aligned bounding-box extents $e_{i}\in\mathbb{R}^{3}_{>0}$.
Graphs and kinematic tree. We build an undirected connection graph $G=(V,E)$ with $V\leftrightarrow P$; an edge indicates geometric contact. A directed kinematic tree $T=(V,E_{T})$, rooted at the base link $b\in V$, orients $G$ and annotates joints.
Joints. For a directed edge $(u\!\to\!v)\in E_{T}$, a joint stores a type $J_{uv}\in\{fixed,revolute,prismatic\}$, a parent-to-child origin $o_{uv}\in\mathbb{R}^{3}$, and if movable, an axis $\mathbf{a}_{uv}\in\mathbb{S}^{2}$ and optionally a pivot $\mathbf{p}_{uv}\in\mathbb{R}^{3}$ when revolute.
III-B Part-Aware 3D Representations
We generate 3D part-level 3D meshes with a 3D foundation model [3] from the input RGB image, and discard meshes with too few vertices or a degenerate spatial spread. For each prospective parent part, we train a continuous SDF [39, 40, 41] $f_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$ on (i) noisy surface points, (ii) near-surface offsets, and (iii) far samples in a bounding AABB.
Afterward, we build a connection graph $G$ with the trained SDF, as shown in the middle left panel of Fig. 1. Given two candidate parts $A$ and $B$ with respective SDFs $f_{\theta_{A}}$ and $f_{\theta_{B}}$, we evaluate mutual distances between their sampled surfaces under $f_{\theta}$. Pairs of parts whose minimum bidirectional distance falls below a tolerance $\epsilon$ are declared in contact, and an undirected edge is added between them.
III-C Kinematic Topology Inference
[…]
Base selection and BFS orientation
We select the base link as any
$b\in\arg\max_{i\in V(G)}\deg_{G}(i),$
where $\deg_{G}(i)$ is the undirected degree in $G$. Starting from $b$, we run BFS on $G$ as a warm start for the MCTS search. During BFS, when a new neighbour $v$ is first reached from an already visited node $u$, we define $u$ as the parent and $v$ as the child, orient the edge as $u\!\to\!v$, and set its origin
$o_{uv}=c_{v}-c_{u}.$ (1)
Any edge whose addition would create a cycle is not inserted into $T$ and is recorded as broken. If $G$ is disconnected, each remaining component is attached to the current tree by adding a virtual edge from its nearest neighbour in the tree, provided the Euclidean distance is at most $d_{\max}$.
III-C1 State, actions, constraints
A search state is $S=(T_{S},V_{S},B_{S})$, where $T_{S}$ is the current partial directed tree, $V_{S}\subseteq V(G)$ the visited-node set, and $B_{S}\subseteq E(G)$ the set of broken undirected edges discovered so far. An action adds a feasible oriented edge $u\!\to\!v$ with $u\in V_{S}$ and $v\notin V_{S}$. To respect discovered symmetries, we form part clusters $\{C_{k}\}$ by the Chamfer distance between segmented meshes. During expansion, we forbid connecting two nodes that belong to the same multi-member cluster to avoid spurious intra-cluster links.
Figure 1: Pipeline of Kinematify for recovering articulated robots from a single RGB image. Step 1: A 3D foundation model generates a segmented mesh of the robot. Step 2: A contact graph is constructed over mesh parts, capturing candidate relations between components. Step 3: MCTS is applied to infer the kinematic tree, resolving ambiguous attachments by leveraging structural priors such as hierarchy and symmetry. Step 4: The DW-CAVL optimization refines joint parameters by preserving near-contact geometry. Bottom row: examples of inferred revolute and prismatic joints with optimized axes.
III-C2 Transition
Applying an action $u\!\to\!v$ updates the edge origin $o_{uv}\leftarrow c_{v}-c_{u}$, provisionally treats the joint as fixed until later typing, inserts $v$ into $V_{S}$, and appends to $B_{S}$ any undirected edge $(v,w)\in E(G)$ with $w\in V_{S}\setminus\{u\}$ that would otherwise form a cycle.
III-C3 Reward
For a completed tree $\tilde{T}$, the terminal reward is a weighted sum of five terms:
$R_{\mathrm{struct}}$.
This term penalises irregular depths, excessive out-degree, and long trees:
$R_{\mathrm{struct}}=\frac{1}{1+\overline{d^{2}}+\overline{(\deg^{+}-3)^{2}}+\lambda\,|E(\tilde{T})|}$, (2)
where $\lambda>0$ is a hyperparameter.
$R_{\mathrm{static}}$.
This term favours centre-of-mass support to reduce gravitational torque about joint frames. Let $v_{i}>0$ denote the estimated volume of part $i$ and $m_{i}=\rho\,v_{i}$ its mass for a density parameter $\rho>0$. With subtree mass $M(i)$ and subtree centre $\mathbf{c}_{\mathrm{sub}}(i)$,
$M(i)$ $=m_{i}+\sum_{j\in\mathrm{ch}(i)}M(j),$ (3)
$\mathbf{c}_{\mathrm{sub}}(i)$ $=\frac{m_{i}\,c_{i}+\sum_{j\in\mathrm{ch}(i)}M(j)\,\mathbf{c}_{\mathrm{sub}}(j)}{M(i)},$ (4)
where $\mathrm{ch}(i)$ is the set of children of $i$ in $\tilde{T}$. Let $\hat{\mathbf{z}}^{-}=[0,\,0,\,-1]^{\top}$, which represents downward unit gravity and $g>0$ the gravitational constant. The total gravitational torque is
$\tau$ $=\sum_{i\neq b}\left\|\big(\mathbf{c}_{\mathrm{sub}}(i)-c_{i}\big)\times\big(M(i)\,g\,\hat{\mathbf{z}}^{-}\big)\right\|_{2},$ (5)
$R_{\mathrm{static}}$ $=\frac{1}{1+\tau/\sigma_{\tau}},$ (6)
with $\sigma_{\tau}>0$ a robust per-assembly normaliser based on MAD scale.
$R_{\mathrm{contact}}$.
We quantify contact strength from SDF-based bidirectional proximity. Let $s(u,v)\in[0,1]$ denote the contact strength of a physical contact on edge $(u\!\to\!v)$. We reward higher average strength:
$R_{\mathrm{contact}}=\frac{1}{|E(\tilde{T})|}\sum_{(u\to v)\in E(\tilde{T})}s(u,v).$ (7)
$R_{\mathrm{sym}}$.
Within each discovered symmetry cluster $C_{k}$ ($|C_{k}|\geq 2$), we prefer equal depths and a shared parent, such as legs attached to the same torso, fingers to the same palm. Let $P_{k}=\{\mathrm{parent}(i):i\in C_{k},\,i\neq b\}$. We define
$S_{k}$ $=\frac{1}{1+\mathrm{Var}\!\big(d(i):i\in C_{k}\big)}\;+\;\Big[1-\frac{|P_{k}|-1}{|C_{k}|-1+\varepsilon}\Big],$ (8)
$R_{\mathrm{sym}}$ $=\mathrm{mean}_{k}\,S_{k}.$
The second term equals $1$ when all parts in $C_{k}$ share the same parent ($|P_{k}|=1$) and decreases linearly as parents diversify. $\varepsilon>0$ avoids division by zero.
$R_{\mathrm{hier}}$.
We discourage children much larger than their parents by estimated volume. With a small $\varepsilon>0$ to avoid divide-by-zero,
$R_{\mathrm{hier}}=\frac{1}{1+\sum_{(u\to v)}\max\!\left\{0,\ \frac{v_{v}}{v_{u}+\varepsilon}-1\right\}}.$ (9)
III-C4 Search
We use Monte Carlo Tree Search (MCTS) with UCT. Each state is $S=(T_{S},V_{S},B_{S})$. From a state $S$, each child $c$ corresponds to applying one feasible action $u\!\to\!v$. Let $Q(c)$ be the cumulative return backed up through child $c$, $N(c)$ its visit count, and $N(\mathrm{parent})$ the visit count of its parent state. With exploration constant $C>0$, selection chooses
$\arg\max_{c}\ \frac{Q(c)}{N(c)}\;+\;C\sqrt{\frac{\ln N(\mathrm{parent})}{N(c)}}.$ (10)
Rollouts greedily complete the tree by repeatedly choosing any available action with the highest immediate score, and the terminal return R(T~)R($\tilde${T}) is backed up along the simulation path. This objective helps resolve symmetric attachments and multi-branch ambiguities at scale. The middle right panel of Fig. 1 shows the kinematics structure after MCTS search.
Input: Connection graph $G=(V,E)$, base $b$, SDFs $\{f_{\theta}\}$, samples $\{P_{v}\}$
Output: Kinematic tree $T=(V,E_{T})$, joints $\{J_{uv}\}$
1
[…]
3 for $t=1$ to $N_{\max}$ do
4 $S\leftarrow S_{0}$; path $\mathcal{P}\leftarrow\emptyset$ ;
5 while $|V_{S}|<|V|$ do
6    if untried edge exists then
7       $a\leftarrow$ pick untried; $S\leftarrow$ Transition(S,a)(S,a); break
8    else
9       $S\leftarrow\arg\max_{c}\frac{Q(c)}{N(c)}+C\sqrt{\ln N(S)/N(c)}$
10    $\mathcal{P}.\text{append}(S)$
11 $\tilde{T}\leftarrow$ greedy rollout from $S$ ;
12 $R\leftarrow$ Reward(T~)($\tilde${T}); backprop $R$ along $\mathcal{P}$ ;
13
[…]
15
16for edge $(u\!\to\!v)\in E_{T}$ do
17   candidates $\leftarrow$ generate from contact stats ;
18 for candidate $(p,u)$ do
19     optimize $\mathcal{J}_{rev}$ or $\mathcal{J}_{pri}$
20  select best class: revolute if $s_{rev}>\zeta s_{pri}$, else prismatic, else fixed ;
21   store $J_{uv}$
return $(T,\{J_{uv}\})$
Algorithm 1 Kinematify
III-D Joint Reasoning
We render orthographic viewsets for the whole assembly and for joint close-ups. For each $(u\!\to\!v)$, we query VLM on the joint viewset and adopt a decision with abstention. If the VLM successfully identifies the joint type, we group joints by child clusters $C_{k}$ and select a representative by majority and correct outliers.
Let $P_{A}=\{\mathbf{a}_{i}\}$ and $P_{B}=\{\mathbf{b}_{j}\}$ be surface samples of two parts $A$ and $B$ in a common frame. We first extract a contact region $\mathcal{C}$ as the union of points on either part whose nearest neighbour on the other part lies within a small threshold. To downweight spurious pairs, each $\mathbf{x}\in\mathcal{C}$ is assigned a weight that decays with its nearest-neighbour distance, using these weights we compute a weighted contact centroid $\boldsymbol{\mu}_{c}$ and a weighted covariance $\boldsymbol{\Sigma}$. The principal direction with the smallest variance provides a hinge-axis estimate $\widehat{\mathbf{u}}_{\mathrm{PCA}}$. In parallel, we obtain a contact normal $\widehat{\mathbf{n}}$ by averaging nearest-point difference vectors across the two surfaces.
We then form a diverse set of revolute candidates $(\mathbf{p},\mathbf{u})$ as follows. Pivots are initialised at the contact centroid ($\mathbf{p}=\boldsymbol{\mu}_{c}$). Axes are drawn from a compact pool that includes $\widehat{\mathbf{u}}_{\mathrm{PCA}}$, the contact normal $\widehat{\mathbf{n}}$, an orthogonal completion $\widehat{\mathbf{u}}_{\perp}$, the principal axes of $\boldsymbol{\Sigma}$, and a few random unit directions. Along each candidate axis $\mathbf{u}$ we place a handful of pivot samples by sliding $\mathbf{p}$ slightly along $\mathbf{u}$ around $\boldsymbol{\mu}_{c}$.
Figure 2: Examples of articulated objects generated by Kinematify. Each row shows different objects across a sequence of joint configurations.
III-D1 Differentiable rigid motions
For a revolute motion $(\mathbf{p},\mathbf{u},\theta)$ with $\|\mathbf{u}\|_{2}=1$,
$\mathbf{y}$ $=\mathbf{p}+\mathbf{R}(\mathbf{u},\theta)\,(\mathbf{x}-\mathbf{p}),$ (11)
$\mathbf{R}(\mathbf{u},\theta)$ $=\cos\theta\,\mathbf{I}+\sin\theta\,[\mathbf{u}]_{\times}+(1-\cos\theta)\,\mathbf{u}\mathbf{u}^{\top},$ (12)
where $\mathbf{u}]_{\times}$ is the $3\times 3$ cross-product matrix. For a prismatic motion with displacement $t\in\mathbb{R}$, $\mathbf{y}=\mathbf{x}+t\,\mathbf{u}$. We parametrise $\mathbf{u}$ by an unconstrained $\mathbf{a}\in\mathbb{R}^{3}$ via $\mathbf{u}=\mathbf{a}/\|\mathbf{a}\|_{2}$, whose Jacobian is
$\frac{\partial\mathbf{u}}{\partial\mathbf{a}}=\frac{1}{\|\mathbf{a}\|_{2}}\left(\mathbf{I}-\frac{\mathbf{a}\mathbf{a}^{\top}}{\|\mathbf{a}\|_{2}^{2}}\right).$ (13)
III-D2 DW-CAVL objective
Let child samples be $\{\mathbf{b}_{i}\}$. A virtual motion parameter $\delta\in\Theta$ denotes either a revolute angle $\theta$ or a prismatic displacement $t$. Let $\Phi_{\delta}$ be the corresponding rigid transform of the child, and define
$s_{0}(\mathbf{x})=f_{\theta}(\mathbf{x}),\qquad s_{\delta}(\mathbf{x})=f_{\theta}\!\big(\Phi_{\delta}(\mathbf{x})\big).$
With hyperparameters volumetric margin $m_{\mathrm{vol}}>0$, logistic sharpness $k>0$, contact band width $\sigma_{c}>0$, and $\varepsilon_{\mathrm{small}}>0$,
$w_{\mathrm{vol}}(s_{0})$ $=\sigma\!\big(-k(s_{0}-m_{\mathrm{vol}})\big),\quad\sigma(z)=\frac{1}{1+e^{-z}},$ (14)
$w_{\mathrm{dist}}(s_{0})$ $=\exp\!\left(-\frac{s_{0}^{2}}{2\sigma_{c}^{2}}\right),\qquad w_{i}=w_{\mathrm{vol}}(s_{0}(\mathbf{b}_{i}))\,w_{\mathrm{dist}}(s_{0}(\mathbf{b}_{i})).$ (15)
The consistency term penalises separation near contact after motion,
$\mathcal{L}_{\mathrm{cons}}(\delta)=\frac{\sum_{i}w_{i}\big[\max\{0,\ s_{\delta}(\mathbf{b}_{i})-m_{\mathrm{vol}}\}\big]^{2}}{\sum_{i}w_{i}+\varepsilon_{\mathrm{small}}}.$ (16)
Using inverse volumetric weights $\tilde{w}_{i}=\sigma\!\big(+k(s_{0}(\mathbf{b}_{i})-m_{\mathrm{vol}})\big)$, the collision term penalises penetration,
$\mathcal{L}_{\mathrm{coll}}(\delta)=\frac{\sum_{i}\tilde{w}_{i}\big[\max\{0,\ -s_{\delta}(\mathbf{b}_{i})-m_{\mathrm{vol}}\}\big]^{2}}{\sum_{i}\tilde{w}_{i}+\varepsilon_{\mathrm{small}}}.$ (17)
For revolute joints we regularise the pivot toward $\boldsymbol{\mu}_{c}$:
$\mathcal{L}_{\mathrm{reg}}=\lambda_{p}\|\mathbf{p}-\boldsymbol{\mu}_{c}\|_{2}^{2},\qquad\lambda_{p}\geq 0.$ (18)
Aggregating over $\delta\in\Theta$ yields
$\mathcal{J}(\mathbf{p},\mathbf{u})=\frac{1}{|\Theta|}\sum_{\delta\in\Theta}\big(\lambda_{c}\mathcal{L}_{\mathrm{cons}}(\delta)+\lambda_{\mathrm{coll}}\mathcal{L}_{\mathrm{coll}}(\delta)\big)+\mathcal{L}_{\mathrm{reg}},$ (19)
with nonnegative weights $\lambda_{c},\lambda_{\mathrm{coll}}$.
III-D3 Candidate selection
We rank many candidates $(\mathbf{p},\mathbf{u})$ on subsampled points using $s(\mathbf{p},\mathbf{u})$, refine the top-K on full points by minimising $\mathcal{J}$, and output scores $1/(1+\mathcal{J})$.
In summary, we infer joint parameters from static geometry and select the candidate with the highest score. The bottom panel in Fig. 1 shows the optimized results for each joint for the input.
TABLE I: Unified baseline comparison across robots sorted by degrees of freedom (DoF). We report Axis Angle Error (°), Axis Position Error (m), and Tree Edit Distance, where lower values are better.
Metric Method Everyday Object UR10e Franka Panda Unitree Go2 Fetch Allegro Unitree H1 Mean
1-8 DoF 6 DoF 7 DoF 12 DoF 13 DoF 16 DoF 19 DoF
$\mathrm{Axis\ Angle\ Error}\!\downarrow$ Articulate Anymesh 35.80 39.67 42.10 53.23 75.60 78.77 79.35 57.79
ArtGS 13.80 25.52 21.30 22.32 53.81 65.59 41.29 34.80
Ours 2.92 5.34 10.42 9.97 23.10 31.39 29.31 16.06
$\mathrm{Axis\ Pos\ Error}\!\downarrow$ Articulate Anymesh 0.19 0.25 0.31 0.41 0.89 0.32 0.74 0.44
ArtGS 0.75 0.97 0.68 1.13 1.93 0.67 1.32 1.06
Ours 0.23 0.27 0.15 0.30 0.71 0.21 0.68 0.36
$\mathrm{TED}\!\downarrow$ AutoURDF 0.27 0.87 1.28 2.21 3.21 4.83 8.13 2.97
Ours 0.13 1.03 0.89 1.97 1.78 1.22 2.23 1.32
Figure 3: Qualitative comparison of articulation recovery on everyday objects across three methods: Kinematify (ours), Articulate Anymesh, and ArtGS. The red line indicates the joint direction.
IV Experiments
We evaluate Kinematify in two settings: (i) everyday articulated objects and (ii) robotic platforms. We follow prior protocols of Articulate Anymesh [22], which use ground-truth segmented meshes from PartNet-Mobility [42, 43] to isolate the impact of 3D segmentation. This enables fair, direct comparisons to baselines that do not accept raw images or texts. We erase the provided kinematic graphs and joint parameters, and reconstruct the mesh. Under this configuration, we compare against the baselines Articulate AnyMesh [22] and ArtGS [19].
For the robotics platforms, we evaluate Kinematify on six commonly used robot models spanning a range of DoF. Because ArtGS and Articulate AnyMesh do not expose explicit kinematic structure, we additionally compare against AutoURDF [13] for kinematics reconstruction performance.
We also conduct experiments on the end-to-end pipelines, starting from RGB images, and evaluate the output quality with ground-truth to quantify the discrepancies.
IV-A Metrics
We report three metrics evaluating both joint parameter and kinematics tree quality.
Axis Angle Error: The angular deviation between the predicted and ground-truth joint-axis directions, and opposite directions are treated as equivalent.
Axis Position Error: The Euclidean distance between predicted and ground-truth pivot positions in the dataset coordinate frame.
Tree Edit Distance: The Tree Edit Distance [44] between the predicted and ground-truth kinematic trees, for instance, the minimal number of node insertions, deletions, or relabelings needed to match the trees.
Figure 4: Demonstration of Kinematify on two high-DoF robots: Unitree Go2 (12 DoF, left) and Unitree H1 (19 DoF, right). For each case, the pipeline starts from a segmented mesh, followed by kinematic tree inference and joint parameter optimization.
IV-B Quantitative results
Everyday Objects. Table I reports a comparison of Kinematify against Articulate Anymesh and ArtGS on the PartNet-Mobility benchmark. Our method achieves the lowest axis angle error among all approaches, indicating superior accuracy in joint orientation estimation. In terms of axis position error, Kinematify also performs competitively, with values close to the best baseline. A qualitative visualization of these comparisons is provided in Fig. 3. Together, these results demonstrate that Kinematify produces precise joint axes and stable pivot placements for everyday objects.
Robots. We further evaluate performance on six robotic systems by measuring both registration quality and body topology reconstruction. As shown in Table I, Kinematify reduces the Tree Edit Distance by a substantial margin on average, reflecting more faithful recovery of kinematic structures. Representative results on Unitree H1 and Go2 are visualized in Fig. 4. These findings highlight the effectiveness of our MCTS-based objective in reasoning about high-DoF, multi-branched kinematic structures, surpassing prior methods in structural consistency.
IV-C End-to-end evaluation
We further evaluate Kinematify in a full end-to-end setting that starts from single RGB images. A part-aware 3D foundation model [3] first produces a segmented mesh. Then, we apply the kinematic reasoning stack unchanged. Because existing baselines do not natively support image to articulated objects at comparable scope, we report absolute performance rather than head to head comparisons.
TABLE II: End-to-end results. Numbers are absolute.
Metric Everyday Objects Fetch Panda
$\mathrm{Axis\ Angle\ Error}\!\downarrow$ 3.78 32.84 14.08
$\mathrm{Axis\ Position\ Error}\ (\mathrm{m})\!\downarrow$ 0.28 0.95 0.22
$\mathrm{TED}\!\downarrow$ 0.67 2.95 1.17
As shown in II, compared to the geometry-only track in Table I, end-to-end errors increase modestly on EO and more noticeably on Fetch and Panda, consistent with their tighter kinematic tolerances.
IV-D Ablation study
We quantify the contribution of each core component by comparing the full method against two ablations: (i) removing the DW-CAVL anchor term so that optimization considers only collision avoidance; and (ii) replacing the MCTS-based kinematic inference with a BFS strategy.
TABLE III: Ablation study on the proposed method, where EO denotes Everyday Object.
Metric Variant EO Fetch Panda
$\mathrm{Axis\ Angle\ Error}\!\downarrow$ w/o MCTS 4.32 28.30 10.92
w/o DW-CAVL 13.94 42.30 29.39
Ours 2.92 23.10 10.42
$\mathrm{Axis\ Pos\ Error}\!\downarrow$ w/o MCTS 0.59 0.97 0.30
w/o DW-CAVL 1.34 1.82 0.97
Ours 0.23 0.71 0.15
$\mathrm{TED}\!\downarrow$ w/o MCTS 0.39 3.32 2.97
w/o DW-CAVL 0.14 1.93 0.98
Ours 0.13 1.78 0.89
As shown in Table III, substituting MCTS with BFS consistently yields larger TED across robots. BFS greedily attaches along local contacts and lacks long range regularization, leading to incorrect parent choices in symmetric substructures and unbalanced trees. In contrast, removing the DW-CAVL anchor does not drastically change the tree but significantly degrades joint parameters. Without an attraction to the contact centroid and near-surface band, the optimizer favors axes that quickly reduce interpenetration yet drift from true pivots. Overall, the full model achieves the best balance.
Figure 5: Kinematify generates URDFs for both the Fetch robot and the drawer, enabling a demonstration of the robot opening the drawer in Isaac Sim and on hardware, with the same models usable for online planning in ROS and MoveIt.
IV-E Real-world robot manipulation
We export the recovered kinematics to URDF and deploy the models in simulation and on a real robot, shown in Fig. 5). From the segmented mesh, Kinematify generates a Fetch URDF and a simple cabinet URDF. For planning, we load both URDFs into a single MoveIt planning scene and derive an SRDF group for the arm.
We use a constraint motion planner [45, 46, 47] as the backend. The task is executed in two stages: (1) reach-to-grasp and (2) constrained pull. In addition to this drawer-opening scenario, we also demonstrate an online planning task of pouring water from a cup into a container, using the same URDF pipeline and motion-planning setup. The same URDFs are used in Isaac Sim and on hardware. In both cases, the arm follows the planned trajectories without collision, showing that the recovered kinematics are physically consistent and directly usable for online planning in ROS and MoveIt.
V Limitations
Two limitations remain. First, our approach assumes that contact regions are reliable articulation cues and that the contact graph is connected. Spurious contacts can mislead the search and leave parts unattached. Second, we restrict the recovered kinematics to a tree structure and thus do not support closed-loop mechanisms, such as four-bar linkages. Future work includes contact reliability estimation and outlier rejection, and extending the solver with loop-closure constraints to handle kinematic loops.
VI Conclusion
We presented Kinematify, an automated pipeline that synthesizes articulated object and robot descriptions directly from RGB images or text prompts. Across both synthetic and real-world inputs, Kinematify improves registration quality and kinematic topology over prior work. We view Kinematify as a step toward open-vocabulary synthesis of high-DoF articulated structures.

### v2

@@
 Such understanding is captured through articulated objects, which are essential for tasks such as physical simulation, motion planning, and policy learning.
-However, creating these models, particularly for complex systems like robots or objects with high degrees of freedom (DoF), remains a significant challenge.
+However, creating these models, particularly for objects with high degrees of freedom (DoF), remains a significant challenge.
 Existing methods typically rely on motion sequences or strong assumptions from hand-curated datasets, which hinders scalability.
-In this paper, we introduce Kinematify, an automated framework that synthesizes articulated objects directly from arbitrary RGB images or text prompts.
+In this paper, we introduce Kinematify, an automated framework that synthesizes articulated objects directly from arbitrary RGB images or textual descriptions.
 Our method addresses two core challenges: (i) inferring kinematic topologies for high-DoF objects and (ii) estimating joint parameters from static geometry.
@@
 We evaluate Kinematify on diverse inputs from both synthetic and real-world environments, demonstrating improvements in registration and kinematic topology accuracy over prior work.
+https://sites.google.com/deemos.com/kinematify
 22footnotetext: Project lead: Qixuan Zhang (zhangqx@deemos.com).11footnotetext: Corresponding authors: Jingyi Yu (yujingyi@shanghaitech.edu.cn), Lan Xu (xulan1@shanghaitech.edu.cn).
 I Introduction
-Enabling robots to effectively interact with and manipulate objects, as well as to model their own articulated structures for self-perception and adaptation, requires a deep understanding of kinematic topologies and joint parameters.
+Enabling robots to effectively interact with objects, as well as to model their own articulated structures for self-perception and adaptation, requires an accurate understanding of kinematic topologies and joint parameters.
 Articulated robot descriptions capture this understanding by encoding geometry, kinematic dependencies, and dynamic constraints in standard formats like the Unified Robot Description Format (URDF) [1].
-These descriptions are essential for core robotic tasks such as manipulation, locomotion, and policy learning.
+These descriptions are essential for robotic tasks such as manipulation, locomotion, and policy learning.
 However, customizing such descriptions for articulated objects remains a significant challenge, demanding substantial manual effort, especially for high degrees of freedom (DoF) systems like humanoids, quadrupeds, and arms.
 This difficulty arises from the labor intensive processes of part-aware 3D modeling, resolving intricate kinematic dependencies, and inferring precise joint parameters [2].
-While recent advances in part-aware 3D generation [3, 4, 5, 6, 7] now enable the on-demand creation of high-quality segmented meshes from text or image prompts, the bottleneck of kinematics inference remains.
+While recent advances in part-aware 3D generation [3, 4, 5, 6, 7] now enable the on-demand creation of high-quality segmented meshes from RGB images or textual descriptions, the bottleneck of kinematics inference remains.
 This challenge has driven robotics researchers to explore high-DoF articulated objects generation approaches.
@@
 While effective, these systems mainly target everyday objects such as laptops, bottles, and drawers, which typically contain only a few moving parts and relatively simple kinematic dependencies.
-In the context of robot self-modelling, related work such as AutoURDF [13] extends this idea to robots, recovering topology and joint types from point-cloud sequences.
+In the context of self-modeling, related work such as AutoURDF [13] extends this idea to robots, recovering topology and joint types from point-cloud sequences.
 However, it presumes motion data and is largely limited to serial-chain structures, whereas high-DoF objects often exhibit multi-branched linkages.
-We introduce Kinematify, a pipeline that synthesizes articulated robot descriptions from RGB images or texts.
-Our method first generates the segmented mesh with a 3D foundation model, then constructs a contact graph and infers a kinematic tree using an MCTS [14, 15, 16] objective that balances hierarchy and structural regularity.
+To address these challenges, we introduce Kinematify, a framework that generates articulated 3D objects from RGB images or texts.
+Kinematify generates the segmented mesh with a part-aware 3D foundation model, such as Rodin [3], then infers the kinematic tree using an MCTS [14, 15, 16] objective that balances hierarchy and structural regularity.
 Subsequently, it estimates joint parameters via DW-CAVL, a novel Distance-Weighted Contact-Aware Virtual Linkage optimization approach.
@@
 Our contributions are:
-  • An automated pipeline for articulated object synthesis.
-Kinematify generates physics-aware, high-DoF articulated objects directly from arbitrary RGB images or text prompts, without requiring motion data or pre-defined articulation priors.
-  • MCTS-based kinematic tree inference.
-We propose a search objective that encodes structural priors like hierarchy and regularity to resolve ambiguous attachments for complex, high-DoF articulated objects.
-  • A geometry-driven joint parameter estimation method.
-Our DW-CAVL algorithm accurately infers revolute and prismatic joint parameters from static geometry alone by optimizing an SDF-based, contact-aware objective under virtual motions.
+  • An open-vocabulary articulated object generation framework.
+Kinematify generates physics-aware articulated objects directly from arbitrary RGB images or textual descriptions, without requiring motion data, training, or pre-defined articulation priors.
+  • A MCTS-based kinematic tree inference approach.
+We propose a search objective that encodes structural priors like hierarchy and regularity to resolve ambiguous attachments for complex, high-DoF articulated objects with multiple branches.
+  • A SDF-driven joint parameter estimation approach.
+The DW-CAVL algorithm accurately infers revolute and prismatic joint parameters from static geometry by optimizing an SDF-based, contact-aware objective under virtual motions.
 II Related Work
@@
 URDFormer [12] trains a transformer to predict a URDF from a single image, relying on a large-scale synthetic dataset of image-URDF pairs.
-Real2Code [10] and Articulate-Anything [11] leverage large language models (LLMs) to generate code-based representations of articulated objects, with the latter using a reinforcement learning loop to refine the model through simulation feedback.
-While powerful in their open-vocabulary capabilities, these methods often produce functionally plausible rather than geometrically precise models and struggle with the complex topologies common in high-DoF objects.
+Real2Code [10] and Articulate-Anything [11] leverage large language models to generate code-based representations of articulated objects, with the latter using a reinforcement learning loop to refine the model through simulation feedback.
+While powerful in their open-vocabulary capabilities, these methods often produce functionally plausible rather than geometrically precise models and struggle with the multi-branch kinematics, which are common in high-DoF objects.
 II-B Robot Self-Modeling
@@
 It demonstrated superior performance in topology inference for serial-chain objects.
-In contrast to robot self-modelling methods that require an observation of the robot, Kinematify operates in a zero-shot physical context, inferring the kinematics topology from the image or text alone.
-These open vocabulary settings make our approach applicable much earlier in the design pipeline, before a physical prototype is built or motion data is available.
-To the best of our knowledge, Kinematify is the first method to construct high-DoF, kinematics-aware articulated objects directly from unlabelled RGB images or texts.
 III Kinematify
-To overcome the limitations of prior self-modelling methods, we introduce Kinematify, a zero-shot framework that constructs high-DoF, kinematics-aware articulated objects directly from unlabelled RGB images or textual descriptions.
+We introduce Kinematify, a framework that generates kinematics-aware articulated objects directly from RGB images or textual descriptions in a zero-shot context.
 III-A Preliminaries
 Assemblies and parts.
-An assembly $\mathcal{A}$ is a finite set of parts $P=\{P_{i}\}_{i=1}^{N}$.
+An assembly $\mathcal{A}$ is a set of parts $P=\{P_{i}\}_{i=1}^{N}$.
 Part $P_{i}$ has a triangulated surface mesh $M_{i}=(V_{i},F_{i})$ with vertices $V_{i}\subset\mathbb{R}^{3}$ and triangular faces $F_{i}\subset\{1,\dots,|V_{i}|\}^{3}$.
@@
 Graphs and kinematic tree.
-We build an undirected connection graph $G=(V,E)$ with $V\leftrightarrow P$; an edge indicates geometric contact.
+We build an undirected connection graph $G=(V,E)$ with $V\leftrightarrow P$, where an edge indicates geometric contact.
 A directed kinematic tree $T=(V,E_{T})$, rooted at the base link $b\in V$, orients $G$ and annotates joints.
@@
 III-B Part-Aware 3D Representations
-We generate 3D part-level 3D meshes with a 3D foundation model [3] from the input RGB image, and discard meshes with too few vertices or a degenerate spatial spread.
+We generate part-level 3D meshes with a part-aware 3D foundation model [3] from the input RGB images or textual description, and discard meshes with too few vertices or a degenerate spatial spread.
 For each prospective parent part, we train a continuous SDF [39, 40, 41] $f_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$ on (i) noisy surface points, (ii) near-surface offsets, and (iii) far samples in a bounding AABB.
@@
 Base selection and BFS orientation
-We select the base link as any
-$b\in\arg\max_{i\in V(G)}\deg_{G}(i),$
-where $\deg_{G}(i)$ is the undirected degree in $G$.
+We choose the base link $b$ as any node in $G$ with the highest undirected degree.
 Starting from $b$, we run BFS on $G$ as a warm start for the MCTS search.
-During BFS, when a new neighbour $v$ is first reached from an already visited node $u$, we define $u$ as the parent and $v$ as the child, orient the edge as $u\!\to\!v$, and set its origin
-$o_{uv}=c_{v}-c_{u}.$ (1)
+During BFS, when a new neighbour $v$ is first reached from an already visited node $u$, we define $u$ as the parent and $v$ as the child, orient the edge as $u\!\to\!v$, and set its origin $o_{uv}\leftarrow c_{v}-c_{u}$.
 Any edge whose addition would create a cycle is not inserted into $T$ and is recorded as broken.
-If $G$ is disconnected, each remaining component is attached to the current tree by adding a virtual edge from its nearest neighbour in the tree, provided the Euclidean distance is at most $d_{\max}$.
+If $G$ is disconnected, each remaining component is attached to the current tree by adding a virtual edge from its nearest neighbor, provided the Euclidean distance is at most $d_{\max}$.
 III-C1 State, actions, constraints
@@
 Step 2: A contact graph is constructed over mesh parts, capturing candidate relations between components.
-Step 3: MCTS is applied to infer the kinematic tree, resolving ambiguous attachments by leveraging structural priors such as hierarchy and symmetry.
-Step 4: The DW-CAVL optimization refines joint parameters by preserving near-contact geometry.
-Bottom row: examples of inferred revolute and prismatic joints with optimized axes.
+Step 3: Infer the kinematic tree using MCTS, resolving ambiguous connections by leveraging structural priors such as hierarchy and symmetry.
+Step 4: Refine joint parameters using the DW-CAVL optimization approach while preserving near-contact geometry.
+Bottom row: Examples of inferred revolute joints with optimized axes.
 III-C2 Transition
@@
 For a completed tree $\tilde{T}$, the terminal reward is a weighted sum of five terms:
-$R_{\mathrm{struct}}$.
-This term penalises irregular depths, excessive out-degree, and long trees:
-$R_{\mathrm{struct}}=\frac{1}{1+\overline{d^{2}}+\overline{(\deg^{+}-3)^{2}}+\lambda\,|E(\tilde{T})|}$, (2)
-where $\lambda>0$ is a hyperparameter.
-$R_{\mathrm{static}}$.
+$R_{\mathrm{struct}}$
+This term penalizes irregular depths, excessive out-degree, and long trees:
+$R_{\mathrm{struct}}=\frac{1}{\overline{d^{2}}+\overline{(\deg^{+}\!-\!k)^{2}}+\lambda\,|E(\tilde{T})|}$, (1)
+where $k$ is the preferred out-degree and $\lambda>0$ is a hyperparameter.
+$R_{\mathrm{static}}$
 This term favours centre-of-mass support to reduce gravitational torque about joint frames.
@@
 where $\mathrm{ch}(i)$ is the set of children of $i$ in $\tilde{T}$.
-Let $\hat{\mathbf{z}}^{-}=[0,\,0,\,-1]^{\top}$, which represents downward unit gravity and $g>0$ the gravitational constant.
+Let $\hat{\mathbf{z}}^{-}=[0,\,0,\,-1]^{\top}$ denote downward unit gravity and $g>0$ the gravitational constant.
 The total gravitational torque is
@@
 with $\sigma_{\tau}>0$ a robust per-assembly normaliser based on MAD scale.
-$R_{\mathrm{contact}}$.
+$R_{\mathrm{contact}}$
 We quantify contact strength from SDF-based bidirectional proximity. Let $s(u,v)\in[0,1]$ denote the contact strength of a physical contact on edge $(u\!\to\!v)$. We reward higher average strength:
 $R_{\mathrm{contact}}=\frac{1}{|E(\tilde{T})|}\sum_{(u\to v)\in E(\tilde{T})}s(u,v).$ (6)
-$R_{\mathrm{sym}}$.
+$R_{\mathrm{sym}}$
 Within each discovered symmetry cluster $C_{k}$ ($|C_{k}|\geq 2$), we prefer equal depths and a shared parent, such as legs attached to the same torso, fingers to the same palm.
@@
 The second term equals $1$ when all parts in $C_{k}$ share the same parent ($|P_{k}|=1$) and decreases linearly as parents diversify. $\varepsilon>0$ avoids division by zero.
-$R_{\mathrm{hier}}$.
+$R_{\mathrm{hier}}$
 We discourage children much larger than their parents by estimated volume. With a small $\varepsilon>0$ to avoid divide-by-zero,
@@
 $w_{\mathrm{dist}}(s_{0})$ $=\exp\!\left(-\frac{s_{0}^{2}}{2\sigma_{c}^{2}}\right),\qquad w_{i}=w_{\mathrm{vol}}(s_{0}(\mathbf{b}_{i}))\,w_{\mathrm{dist}}(s_{0}(\mathbf{b}_{i})).$ (14)
-The consistency term penalises separation near contact after motion,
+The consistency term penalizes separation near contact after motion,
 $\mathcal{L}_{\mathrm{cons}}(\delta)=\frac{\sum_{i}w_{i}\big[\max\{0,\ s_{\delta}(\mathbf{b}_{i})-m_{\mathrm{vol}}\}\big]^{2}}{\sum_{i}w_{i}+\varepsilon_{\mathrm{small}}}.$ (15)
@@
 Overall, the full model achieves the best balance.
-Figure 5: Kinematify generates URDFs for both the Fetch robot and the drawer, enabling a demonstration of the robot opening the drawer in Isaac Sim and on hardware, with the same models usable for online planning in ROS and MoveIt.
+Figure 5: Kinematify generates URDFs for both the Fetch robot and the drawer, enabling a demonstration of the robot opening the drawer in Isaac Sim and transfer to real, with the same models usable for online planning with MoveIt.
 IV-E Real-world robot manipulation
@@
 In both cases, the arm follows the planned trajectories without collision, showing that the recovered kinematics are physically consistent and directly usable for online planning in ROS and MoveIt.
-V Limitations
-Two limitations remain.
-First, our approach assumes that contact regions are reliable articulation cues and that the contact graph is connected.
-Spurious contacts can mislead the search and leave parts unattached.
-Second, we restrict the recovered kinematics to a tree structure and thus do not support closed-loop mechanisms, such as four-bar linkages.
-Future work includes contact reliability estimation and outlier rejection, and extending the solver with loop-closure constraints to handle kinematic loops.
+V Discussion and Limitations
+This work addresses the open-vocabulary synthesis of articulated objects from static inputs.
+However, accurate part segmentation and a mechanically faithful contact graph remain assumptions.
+Spurious seams or missed contacts can mislead the search or leave subassemblies unattached.
+A practical extension is to estimate contact reliability and jointly refine segmentation and contacts.
+We also find that decorative geometry near interfaces can introduce false contacts that bias the search.
+In such cases, reweighting the MCTS objective to emphasize the symmetry prior, i.e., increasing the weight on $R_{\mathrm{sym}}$ relative to $R_{\mathrm{contact}}$, helps suppress attachments to decorative elements and favors globally coherent parents within symmetric clusters.
+An adaptive schedule that upweights $R_{\mathrm{sym}}$ when cluster cardinality is high or contact variance is large is a simple and effective mitigation.
+Learning these weights from validation performance is a complementary direction.
 VI Conclusion
-We presented Kinematify, an automated pipeline that synthesizes articulated object and robot descriptions directly from RGB images or text prompts.
-Across both synthetic and real-world inputs, Kinematify improves registration quality and kinematic topology over prior work.
+We presented Kinematify, an automated pipeline that synthesizes articulated object and robot descriptions directly from RGB images or textual descriptions.
+Across everyday objects and robot platforms, Kinematify improves joint accuracy and kinematic tree fidelity over prior work.
 We view Kinematify as a step toward open-vocabulary synthesis of high-DoF articulated structures.